OpenStudy (anonymous):

Use the factor theorem to prove that x+c is a factor of x^n + c^n if n >_ 1 (greater than or equal to sign: >_ ) is an odd integer.

OpenStudy (ash2326):

We need to prove that x+c is a factor of \[x^n+c^n\ if\ n\ge 1\] If x+c is a factor of \(f(x)=(x^n+c^n)\) then f(-c)=0 let's put x=-c \[(-c)^n+c^n\] We know that if n=even, \((-c)^n=c^n\) and if n is odd \(=>(-c)^n=-c^n\) so if n is odd \[f(-c)=-c^n+c^n=0\] Hence x+c is a factor of \(x^n+c^n\) for n=odd

OpenStudy (ash2326):

@ap2332 did you understand?

OpenStudy (anonymous):

yes, thank you. this is very helpful